22 - Celsius Corp. is conducting a capital budgeting analysis to decide whether to invest in...

22 -

Celsius Corp. is conducting a capital budgeting analysis to decide whether to invest in a new project which has an expected life of 5 years. The following information is available:

  • The installed cost of the new equipment is $360,000. Installation will cost an additional $40,000 The equipment will be depreciated using 5-year MACRS depreciation over the 5-year life of the project.
  • An initial NOWC investment equal to 10 percent of year 1 sales will also be required.
  • The equipment is expected to have a salvage value of $80,000 at the end of the project's life.
  • Forecasted sales in year 1 is $600,000 with a gross margin (excluding depreciation) of 40%. Sales is expected to increase 5% per year over the 5-year life of the project.
  • Additional NOWC investments equal to 10 percent of the expected increase in sales will also be required each year.
  • Interest expense from a loan used to finance the project will result in annual interest payments of $28,000 each year over the 5-year life of the project.
  • Celsius has a 21% corporate tax rate.

What is the expected net cash flow after tax (CFAT) in year 1 for the project?

Select one:

a. $203,400

b. $238,200

c. $276,080

d. $298,200

e. none of the above
Answer #1

rate positively .. let me know if you need any clarification..

Computation of initial investment
cost of the new equipment 360000
Installation cost 40000
Total cost of equipment= 400000
1st year depreciation @20% 80000
i sales = 600000
ii=i*40% Gross margin =40% 240000
iii Depreciation = 80000
iv=ii-iii Profit before tax = 160000
v=iv*21% Tax @ 21% 33600
vi=iv-v Profit after tax = 126400
vii=vi+iii Operating cash flow = 206400
viii Working capital =
600000*5%*10% 3000
ix=vii-viii Cash flow after tax = 203400
therefore answer = 203400